OpenStudy (anonymous):

Can someone teach me how to solve a system of equations with three variables through elimination? I am learning this online, but for some reason, I read it over and over and it doesn't make sense.

OpenStudy (anonymous):

If you want to view my lesson, it is here: http://broward.flvs.net/webdav/educator_algebra2_v10/module02/02_04a_a.htm

OpenStudy (anonymous):

http://www.ehow.com/how_8650416_solve-systems-equations-activity.html scroll down, it will teach u the steps and how to.

OpenStudy (anonymous):

that is for a equations with 2 variables, @seiga

OpenStudy (anonymous):

haha like a question like this x + y - 3z = -8 2x + 2y + z = 12 3x + y - z = -2

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

haha a person is help me with one right now, you can come join in on our question if you want